# HN-MVTS: HyperNetwork-based Multivariate Time Series Forecasting

### v1

@@ Abstract
Accurate forecasting of multivariate time series data remains a formidable challenge, particularly due to the growing complexity of temporal dependencies in real-world scenarios. While neural network-based models have achieved notable success in this domain, complex channel-dependent models often suffer from performance degradation compared to channel-independent models that do not consider the relationship between components but provide high robustness due to small capacity. In this work, we propose HN-MVTS, a novel architecture that integrates a hypernetwork-based generative prior with an arbitrary neural network forecasting model. The input of this hypernetwork is a learnable embedding matrix of time series components. To restrict the number of new parameters, the hypernetwork learns to generate the weights of the last layer of the target forecasting networks, serving as a data-adaptive regularizer that improves generalization and long-range predictive accuracy. The hypernetwork is used only during the training, so it does not increase the inference time compared to the base forecasting model. Extensive experiments on eight benchmark datasets demonstrate that application of HN-MVTS to the state-of-the-art models (DLinear, PatchTST, TSMixer, etc.) typically improves their performance. Our findings suggest that hypernetwork-driven parameterization offers a promising direction for enhancing existing forecasting techniques in complex scenarios.
@@ Introduction
Multivariate time series (MVTS) forecasting (Kostromina et al. 2025; Mendis, Wickramasinghe, and Marasinghe 2024) is critical to many real-world applications, including energy demand prediction, traffic management, financial forecasting, weather modeling, and health monitoring. In such scenarios, multiple interrelated time series, often called channels or components, must be predicted simultaneously. Effective forecasting, therefore, depends not only on modeling temporal patterns within each series but also on capturing correlations across components.
@@
Recent advancements in deep learning have led to powerful forecasting models, ranging from linear baselines (Zeng et al. 2023) to architectures based on convolutional networks (Luo and Wang 2024), multilayer perceptrons (MLPs) (Chen et al. 2023), and Transformers (Nie et al. 2023). Surprisingly, recent work has shown that channel-independent (CI) models, those that treat each time series channel separately, often outperform channel-dependent (CD) approaches that attempt to jointly model all channels (Han, Ye, and Zhan 2024; Peiwen and Changsheng 2023). CI models reduce model complexity, but sacrifice the ability to leverage valuable inter-channel relationships, potentially limiting their accuracy in highly correlated settings.
@@
On the other hand, CD models explicitly exploit inter-channel dependencies. Indeed, the improvement over baseline is often achieved by advanced modeling of temporal (Nie et al. 2023) and channel (Liu et al. 2024b) similarities, or both (Wu et al. 2023). However, CD models are known to be often less scalable due to their need for larger datasets (Peiwen and Changsheng 2023). Designing models that can balance the robustness of CI approaches with the expressiveness of CD strategies remains a key open challenge in MVTS forecasting (Wang et al. 2023; Zhao and Shen 2024). The difficulty lies in balancing leveraging inter-channel correlations and maintaining the robustness and scalability across diverse datasets. For example, a local model with individual parameters for each channel leads to increased parameters and does not model the channels’ similarities (Montero-Manso and Hyndman 2021).
@@
To address this, we propose HN-MVTS (Fig. 1), a novel architecture that bridges the gap between channel-independent and channel-dependent modeling by using a hypernetwork to generate channel-specific forecasting parameters. Specifically, a small MLP-based hypernetwork receives a learnable embedding for each channel and outputs the weights of the final prediction layer for that channel. This setup allows similar channels to share information through their embeddings while preserving the modularity and efficiency of CI models. A key strength of HN-MVTS is its architectural flexibility: it can be seamlessly integrated with a wide range of neural network-based forecasting models, including linear baselines, MLPs, CNNs, and Transformer variants. By modifying only the final prediction layer, HN-MVTS enhances the forecasting performance of existing architectures with minimal overhead in training time and the number of learnable parameters and without any impact on the inference time. We demonstrate the effectiveness of HN-MVTS through comprehensive experiments on multiple benchmark datasets. Our approach typically substantially improves the performance of several state-of-the-art models, such as DLinear, PatchTST, and TSMixer, on multiple widely-used datasets (ECL, ETTh, etc.), highlighting its robustness, adaptability, and ability to generalize across diverse settings.
@@
Models. Nowadays, the dominant architecture in multivariate time series forecasting is the Transformer. Modern papers illustrate an expanding design space where model efficacy increasingly depends on judicious architectural alignment with time series’ unique characteristics rather than computational complexity alone (Su et al. 2025; Wang et al. 2024). The Inverse Transformer (iTransformer) (Liu et al. 2024b), processes entire time points as variate tokens rather than temporal tokens, enabling better multivariate correlation capture while retaining Transformer components. It addresses key limitations of conventional temporal tokenization, particularly when handling large lookback windows and diverse variate relationships. PatchTST (Nie et al. 2023) demonstrates how patching strategies from computer vision can adapt Transformers for time series through subseries-level embeddings and channel independence. However, recent works have introduced various models that challenge traditional architectures by focusing on improved temporal and cross-variable dependencies. For example, the DLinear architecture (Zeng et al. 2023) emerged as a compelling alternative by decomposing series into trend and seasonality components through linear layers, achieving superior performance over complex Transformer variants like FEDformer (Zhou et al. 2022) on key benchmarks while maintaining computational efficiency. This simplicity-first approach sparked reevaluations of model design priorities, with ModernTCN (Luo and Wang 2024) further advancing non-Transformer paradigms through modernized temporal convolutional networks that expand effective receptive fields while maintaining computational efficiency. TSMixer (Chen et al. 2023), TimesMixer (Wang et al. ) and HDMixer (Huang et al. 2024) exemplify this through an all-MLP (multi-layer perceptron) architecture that integrates historical data, future known inputs, and static covariates using conditional feature mixing layers, demonstrating competitive performance against Transformers on electricity datasets. Finally, there exists techniques based on graph neural networks, e.g., graph inter-series models, such as MSGNet (Cai et al. 2024a) and FourierGNN (Yi et al. 2023), or ForecastGrapher (Cai et al. 2024b) that treats each series as a node and learns a graph, but there scalability is usually rather poor.
@@
Channel-dependent vs channel-independent models. MVTS forecasting strategies are broadly categorized into CI and CD (Montero-Manso and Hyndman 2021). The CI mode (Han, Ye, and Zhan 2024), which includes both global (single model applied to all univariate series) and local (dedicated models per series) options, treats each time series channel as independent, ignoring cross-channel correlations. In contrast, the CD strategy employs a single model that takes all components as input and forecasts all components simultaneously, explicitly modeling inter-channel dependencies (Hertel et al. 2023). The CI-CD performance gap stems from fundamental trade-offs between model capacity and robustness. CD methods theoretically possess the higher capacity for modeling complex channel interactions but struggle with real-world challenges like limited training data. CI strategies overcome these issues by using larger effective datasets (through channel-wise replication) and avoiding error propagation from correlated channels. Hence, empirical studies (Hertel et al. 2023) demonstrate that CI methods outperform CD approaches across diverse domains, including energy load forecasting, financial data, and epidemiological modeling. (Han, Ye, and Zhan 2024) benchmarks global univariate versus multivariate modeling for linear, autoregressive, MLP, gradient boosting, and transformer-based models. They show that CI (univariate) models generally significantly outperforming multivariate ones, though most transformer-based MVTS forecasting models are trained using the multivariate strategy. Global CI methods achieve parameter efficiency comparable to local models while maintaining competitive accuracy (Montero-Manso and Hyndman 2021). Indeed, the top linear model (DLinear) (Zeng et al. 2023) trained using the global univariate approach outperforms multivariate-trained transformers. Similarly, (Hertel et al. 2023) reports that global CI Transformers significantly reduce forecasting errors compared to CD counterparts in smart grid applications. The success of CI highlights the importance of robustness in real-world forecasting (Ashouri and Phoa 2022). Nevertheless, CD approaches retain promise for domains with strong inter-channel correlations, provided they incorporate mechanisms like temporal clustering or residual regularization (Han, Ye, and Zhan 2024). For example, DUET (Qiu et al. 2024) introduces dual temporal-channel clustering to enhance CD models, showing that structured dependency modeling combined with robustness mechanisms can match CI performance in scenarios with clear channel groupings. Above-mentioned PatchTST (Nie et al. 2023) processes time series into patched “words” using CI Transformers to preserve univariate patterns while enabling cross-variate attention. (Chen et al. 2024) observes the boost in forecasting performance for multivariate models for correlated time series. These developments underscore the need for context-aware strategy selection, balancing the CI-CD trade-off against dataset characteristics and operational constraints. Similar challenges are addressed in COSCI-GAN (Seyfi, Rajotte, and Ng 2022) for another task, generative modeling, by coordinating multiple single-channel generators from a common noise source and using a central discriminator to enforce inter-channel relationship, but does not perform embedding-based parameter generation approach.
@@
Hypernetworks. Hypernetworks, neural networks that generate weights for another neural network (the target network), have emerged as a powerful approach for various machine learning tasks (Ha, Dai, and Le 2016), including time series forecasting. The foundational work (Duan et al. 2023) introduced Hyper Time Series Forecasting, which jointly learns time-varying distributions and corresponding forecasting models, enabling more accurate predictions when underlying data distributions change over time. Recent advances include HyperGPA (Lee et al. 2022), a hypernetwork that generates optimal model parameters for each time period through computational graph structures. It demonstrates hypernetworks’ effectiveness in forecasting scenarios where future data patterns differ from historical observations. Finally, LPCNet employs a hypernetwork-based framework that dynamically updates neural network parameters in response to changing input distributions (Liu et al. 2024a). This adaptive parameter generation automatically enables the model to adjust to varying temporal patterns with acceptable computational complexity. These advancements highlight hypernetworks’ ability to model complex temporal dependencies while adapting to evolving data distributions.
@@
Another research direction is the usage of a hypernetwork for meta-learning to handle multiple tasks. Hyper-Trees (März and Rasul 2024) generate parameters that adapt to local characteristics when applied to individual series, enabling effective handling of heterogeneous time series. Recent work (Staněk 2025) explored the hypernetwork’s meta-learning architectures capable of constructing optimal parametric models for families of similar data-generating processes, achieving competitive performance in forecasting challenges. The HyperTime framework (Fons et al. 2022) provides an accurate and resolution-independent encoding of time series data, facilitating accurate reconstruction and interpolation. It was particularly effective for data augmentation, achieving competitive performance against state-of-the-art time series augmentation methods.
@@
Thus, hypernetworks show promise for limited observations where nonparametric methods fail. However, their applications are still restricted to managing non-stationary time series, data augmentation, or training when multiple realizations of similar but not identical time series are available. To the best of our knowledge, hypernetworks have not been applied to improve the performance metrics, such as MSE (mean squared error), of the state-of-the-art models for MVTS forecasting. This paper fills this gap by using hypernetworks to take advantage of both CI and CD approaches.
@@
Let the history be , where is the number of channels, and is the length of the lookback window. The multivariate time series forecasting problem is to predict , where is the prediction horizon.
@@
In this paper, we focus on neural network models with layers, and assume that its last (-th) layer linearly transforms an input hidden state , where with dimensionality into the final output using the matrix product for each channel :
@@
In this paper, we implement a hypernetwork (3) as a conventional small MLP that that takes a learnable embedding vector for each time series component (treating the number of channels as the batch size) and outputs the corresponding weights for the final prediction layer (2) of an arbitrary base forecasting model. In the simplest case of an MLP without hidden layers, the weights of the base model are obtained using a linear transform:
@@
The central contribution of our method lies in using a hypernetwork to generate part of the parameters of a forecasting model in a CD yet scalable and data-adaptive manner. In traditional multivariate forecasting models, all output channels typically share the same output projection layer (CD), or each channel has its separate model (CI). Our approach aims to combine both benefits by introducing a shared neural forecasting model, while tailoring its final layer to each channel via a learnable hypernetwork. The difference between CD, CI, and our hypernetwork-based approach (HN-MVTS) is shown in Fig. 1. The former is the most common model, with a single transform of input data into outputs. The CI approach learns separate models , and processes each channel independently. Our HN-MVTS can save the advantages of both types of models. If some components and of time series are similar, their embeddings will be close to each other , and, hence, the training data for the component will have more influence to learn the weights for the component, and vice versa. In the ideal case of clustered embeddings , a single (global) model (Montero-Manso and Hyndman 2021) will be trained for the union of these components. In contrast, if embeddings of and components are significantly different, a CI mode is activated, in which the forecasting parameters are trained independently. This mechanism allows the model to adaptively interpolate between CD and CI behavior based on the learned embedding, without manual architecture changes or channel grouping.
@@
Some examples of adding our HN-MVTS to various architectures are shown in Fig. 2. The above-mentioned motivation of mixing CD and CI models is especially clear for a forecasting model with one linear layer (left part of this figure), where the columns of the weight matrix are learned based on the similarity between embeddings of time series components. However, our approach can be easily applied to MLP (Fig. 2b) or more complex architectures such as transformers (Fig. 2c) as only the weights of the last prediction layer are defined by the hypernetwork. It may be used for all layers, but the number of trained parameters will be much greater, making it difficult to train the model. A practitioner can control the number of excess parameters by managing the complexity of the hypernetwork’s MLP.
@@
In all experiments, we use the following straightforward implementation of HN-MVTS. First, we initialize embeddings from the Pearson product-moment correlation coefficients across channels (Nguyen-Thai et al. 2024), projected onto the principal components of dimension computed for the training-only splits. Our preliminarily experiments show that alternative (random) initialization leads to slightly higher MSE and/or increased training time. Next, we attach a simple MLP (4) and learn its weights and embeddings from given data together with the base model (1) by minimizing MSE. Our simplest implementation (4) adds only parameters to learn (with addition of if embeddings are learnable). As a result, the number of parameters can be even less than that of the CI approach, which applies the base model separately with its parameters for each channel. As the number of new learnable weights is reasonably small, the training time should be only slightly higher compared to the base forecasting model. Moreover, as the trained hypernetwork outputs the fixed weights that do not depend on the concrete input time series, we can compute them only once after the training procedure. Hence, we remove a hypernetwork, and copy weights (4) into the last linear layer of the base model (2). As a result, the inference time won’t be affected (though deployment overheads, e.g., I/O, memory layout, may still cause minor variation), making our HN-MVTS as fast as the base model.
@@
Thus, the proposed approach uses a learnable component embedding combined with a hypernetwork to produce channel-specific final-layer weights as a data-adaptive regularizer, integrate seamlessly with diverse backbone architectures, and preserve inference cost. These points with empirical gains discussed in the next section are a substantive contribution beyond straightforward reapplication of existing hypernetwork’s ideas.
@@
We evaluated the proposed HN-MVTS with contemporary multivariate time series forecasting models of different architectures: 1) Linear model: DLinear (Zeng et al. 2023), 2) MLP-based model: TSMixer (Chen et al. 2023), 3) Convolutional neural network-based model: ModernTCN (Luo and Wang 2024), and 4) Transformer-based models: PatchTST (Nie et al. 2023) and inverse Transformer (Liu et al. 2024b). The source code for all models was taken from original repositories of their authors.
@@
We leverage eight publicly available open datasets that are conventionally used in the MVTS literature (Shao et al. 2025), covering energy, traffic, and weather forecasting applications: ECL (UCI Electricity Consuming Load) with electricity consumption of hunderds of points/clients; 15-minute-level datasets (ETTm1, ETTm2) from the ETT (Electricity Transformer Temperature) (Zhou et al. 2021) with one oil and six load features of electricity transformers from 2016 to 2018; Weather (Max-Planck-Institute Weather Dataset for Long-term Time Series Forecasting) with weather indicators recorded every 10 min for 2020, and Traffic Flow from Performance Measurement System (PEMS03, PEMS04, PEMS07, PEMS08) collected by California Transportation Agencies (CalTrans) every 30 seconds and aggregated to 5 minutes. PEMS03, PEMS07 and PEMS08 contain three months of traffic flow from 358, 883, and 170 sensors, respectively, while PEMS04 includes two months of statistics from 307 sensors.
Each dataset exhibits diverse temporal and structural properties, allowing us to assess the robustness and generalizability of our method. Their statistics are summarized in Table 1. We split time series for train, validation, and test using ratio. For ETTh and ETTm datasets, following the standard testing protocols, we use a split and take only the first and timesteps, respectively. The validation set was used to obtain the best checkpoint of the model, and the results on the test set are reported.

### v2

@@ Abstract
Accurate forecasting of multivariate time series data remains a formidable challenge, particularly due to the growing complexity of temporal dependencies in real-world scenarios. While neural network-based models have achieved notable success in this domain, complex channel-dependent models often suffer from performance degradation compared to channel-independent models that do not consider the relationship between components but provide high robustness due to small capacity. In this work, we propose HN-MVTS, a novel architecture that integrates the hypernetwork-based generative prior with an arbitrary neural network forecasting model. The input of this hypernetwork is a learnable embedding matrix of time series components. To restrict the number of new parameters, the hypernetwork learns to generate the weights of the last layer of the target forecasting networks, serving as a data-adaptive regularizer that improves generalization and long-range predictive accuracy. The hypernetwork is only used during training, so it does not increase the inference time compared to the base forecasting model. Extensive experiments on eight benchmark datasets demonstrate that application of HN-MVTS to the state-of-the-art models (DLinear, PatchTST, TSMixer, etc.) typically improves their performance. Our findings suggest that hypernetwork-driven parameterization offers a promising direction for enhancing existing forecasting techniques in complex scenarios.
@@ Introduction
Multivariate time series (MVTS) forecasting (Kostromina et al. 2025; Mendis, Wickramasinghe, and Marasinghe 2024) is critical to many real-world applications, including energy demand prediction, traffic management, financial forecasting, weather modeling, and health monitoring. In such scenarios, multiple interrelated time series, often referred to as channels or components, must be predicted simultaneously. Effective forecasting, therefore, depends not only on modeling temporal patterns within each series but also on capturing correlations across components.
@@
Advancements in deep learning have led to powerful forecasting models, ranging from linear baselines (Zeng et al. 2023) to architectures based on convolutional networks (Luo and Wang 2024), multilayer perceptrons (MLPs) (Chen et al. 2023), and Transformers (Nie et al. 2023). Surprisingly, recent work has shown that channel-independent (CI) models, which treat each time series channel separately, often outperform channel-dependent (CD) approaches that attempt to model all channels jointly (Han, Ye, and Zhan 2024; Peiwen and Changsheng 2023). CI models reduce model complexity, but sacrifice the ability to leverage valuable inter-channel relationships, potentially limiting their accuracy in highly correlated settings.
@@
On the other hand, CD models explicitly exploit inter-channel dependencies. Indeed, the improvement over baseline is often achieved by advanced modeling of temporal (Nie et al. 2023) and channel (Liu et al. 2024b) similarities, or both (Wu et al. 2023). However, CD models are often known to be less scalable due to their requirement for larger datasets (Peiwen and Changsheng 2023). Designing models that can balance the robustness of CI approaches with the expressiveness of CD strategies remains a key open challenge in MVTS forecasting (Wang et al. 2025; Zhao and Shen 2024). The difficulty lies in balancing the leverage of inter-channel correlations and maintaining robustness and scalability across diverse datasets. For example, a local model with individual parameters for each channel results in increased parameters and does not capture the channels’ similarities (Montero-Manso and Hyndman 2021).
@@
To address this, we propose HN-MVTS (Fig. 1), a novel architecture that bridges the gap between channel-independent and channel-dependent modeling by using a hypernetwork to generate channel-specific forecasting parameters. Specifically, a small MLP-based hypernetwork receives a learnable embedding for each channel and outputs the weights of the final prediction layer for that channel. This setup allows similar channels to share information through their embeddings while preserving the modularity and efficiency of CI models. A key strength of HN-MVTS is its architectural flexibility: it can be seamlessly integrated with a wide range of neural network-based forecasting models, including linear baselines, MLPs, ConvNets, and Transformer variants. By modifying only the final prediction layer, HN-MVTS enhances the forecasting performance of existing architectures with minimal overhead in training time and the number of learnable parameters and without any impact on the inference time. We demonstrate the effectiveness of HN-MVTS through comprehensive experiments on multiple benchmark datasets. Our approach typically substantially improves the performance of several state-of-the-art models, such as DLinear, PatchTST, and TSMixer, on multiple widely used datasets (e.g., ECL, ETTh), highlighting its robustness, adaptability, and ability to generalize across diverse settings.
@@
Models. Nowadays, the dominant architecture in MVTS forecasting is the Transformer. Modern papers illustrate an expanding design space where model efficacy increasingly depends on architectural alignment with the unique characteristics of time series (Su et al. 2025; Wang et al. 2024b). Inverse Transformer (iTransformer) (Liu et al. 2024b) processes entire time points as variate tokens rather than temporal tokens, enabling better capture of multivariate correlations while retaining Transformer components. It addresses key limitations of conventional temporal tokenization, particularly when handling large lookback windows and diverse variate relationships. PatchTST (Nie et al. 2023) demonstrates how patching from computer vision can adapt Transformers for time series through subseries-level embeddings and channel independence.
@@
However, recent works have introduced various models that challenge traditional architectures by focusing on improved temporal and cross-variable dependencies. For example, the DLinear architecture (Zeng et al. 2023) has emerged as a compelling alternative by decomposing series into trend and seasonality components through linear layers, achieving superior performance over complex Transformer variants, such as FEDformer (Zhou et al. 2022), on key benchmarks while maintaining computational efficiency. This simplicity-first approach sparked reevaluations of model design priorities, with ModernTCN (Luo and Wang 2024) further advancing non-Transformer paradigms through modernized temporal convolutional networks that expand effective receptive fields while maintaining computational efficiency. TSMixer (Chen et al. 2023), TimesMixer (Wang et al. 2024a), and HDMixer (Huang et al. 2024) exemplify this through an all-MLP (multi-layer perceptron) architecture that integrates historical data, future known inputs, and static covariates using conditional feature mixing layers, demonstrating competitive performance against Transformers.
Finally, techniques based on graph neural networks (Makarov et al. 2022), such as graph inter-series models like MSGNet (Cai et al. 2024a) and FourierGNN (Yi et al. 2023), or ForecastGrapher (Cai et al. 2024b), treat each series as a node and learn a graph; however, their scalability is usually relatively poor.
@@
Channel-dependent vs channel-independent models. MVTS forecasting strategies are broadly categorized into CI and CD (Montero-Manso and Hyndman 2021). The CI mode (Han, Ye, and Zhan 2024), which includes both global (single model applied to all univariate series) and local (dedicated models per series) options, treats each time series channel as independent, ignoring cross-channel correlations. In contrast, the CD strategy employs a single model that takes all components as input and forecasts all components simultaneously, explicitly modeling inter-channel dependencies (Hertel et al. 2023). The CI-CD performance gap stems from fundamental trade-offs between model capacity and robustness. CD methods theoretically possess a higher capacity for modeling complex channel interactions, but struggle with real-world challenges, such as limited training data. CI strategies overcome these issues by using larger effective datasets (through channel-wise replication) and avoiding error propagation from correlated channels. Hence, empirical studies (Hertel et al. 2023) demonstrate that CI methods outperform CD approaches across various domains, including energy load forecasting, financial data analysis, and epidemiological modeling. (Han, Ye, and Zhan 2024) benchmarks global univariate versus multivariate modeling for linear, autoregressive, MLP, gradient boosting, and transformer-based models. They show that CI (univariate) models generally significantly outperform multivariate ones, though most transformer-based MVTS forecasting models are trained using the multivariate strategy. Global CI methods achieve parameter efficiency comparable to local models while maintaining competitive accuracy (Montero-Manso and Hyndman 2021). Indeed, the top linear model (DLinear) (Zeng et al. 2023) trained using the global univariate approach outperforms multivariate-trained transformers. Similarly, (Hertel et al. 2023) reports that global CI Transformers significantly reduce forecasting errors compared to CD counterparts in smart grid applications. The success of CI highlights the importance of robustness in real-world forecasting (Ashouri and Phoa 2022). Nevertheless, CD approaches retain promise for domains with strong inter-channel correlations, provided they incorporate mechanisms like temporal clustering or residual regularization (Han, Ye, and Zhan 2024). For example, DUET (Qiu et al. 2024) introduces dual temporal-channel clustering to enhance CD models, showing that structured dependency modeling combined with robustness mechanisms can match CI performance in scenarios with clear channel groupings. PatchTST (Nie et al. 2023) processes time series into patched “words” using CI Transformers to preserve univariate patterns while enabling cross-variate attention. (Chen et al. 2024) observes the boost in forecasting performance for multivariate models for correlated time series. These developments underscore the need for context-aware strategy selection, balancing the CI-CD trade-off against dataset characteristics and operational constraints. Similar challenges are addressed in COSCI-GAN (Seyfi, Rajotte, and Ng 2022) for another task, generative modeling, by coordinating single-channel generators from a common noise source and utilizing a central discriminator to enforce inter-channel relationships. However, it does not employ an embedding-based parameter generation approach.
@@
Hypernetworks. Hypernetworks, neural networks that generate weights for another neural network (the target network), have emerged as a powerful approach for various machine learning tasks (Ha, Dai, and Le 2016), including time series forecasting. The foundational work (Duan et al. 2023) introduced Hyper Time Series Forecasting, which jointly learns time-varying distributions and corresponding forecasting models, enabling more accurate predictions when underlying data distributions change over time. Recent advances include HyperGPA (Lee et al. 2022), a hypernetwork that generates optimal model parameters for each time period through computational graph structures. It demonstrates the effectiveness in forecasting scenarios where future data patterns differ from historical observations. Finally, LPCNet employs a hypernetwork-based framework that dynamically updates neural network parameters in response to changing input distributions (Liu et al. 2024a). This adaptive parameter generation automatically enables the model to adjust to varying temporal patterns with acceptable computational complexity. These advancements highlight the ability of hypernetworks to model complex temporal dependencies while adapting to evolving data distributions.
@@
Another research direction is the usage of a hypernetwork for meta-learning to handle multiple tasks. Hyper-Trees (März and Rasul 2024) generate parameters that adapt to local characteristics when applied to individual series, enabling effective handling of heterogeneous time series. Recent work (Staněk 2025) explored the hypernetwork’s meta-learning architectures, which are capable of constructing optimal parametric models for families of similar data-generating processes, achieving competitive performance in forecasting challenges. The HyperTime framework (Fons et al. 2022) provides an accurate and resolution-independent encoding of time series data, facilitating accurate reconstruction and interpolation. It was effective for data augmentation, achieving competitive performance against state-of-the-art time series augmentation methods.
@@
Thus, hypernetworks show promise for limited observations where nonparametric methods fail. However, their applications are still restricted to managing non-stationary time series, data augmentation, or training when multiple realizations of similar but not identical time series are available. To the best of our knowledge, they have not been applied to improve the performance metrics, such as mean squared error (MSE), of state-of-the-art models for MVTS forecasting. This paper fills this gap by utilizing hypernetworks to leverage the benefits of both CI and CD approaches.
@@
Let the history be , where is the number of channels, and is the length of the lookback window. The multivariate time series forecasting task is to predict , where is the prediction horizon.
@@
In this paper, we focus on neural network models with layers, and assume that its last (-th) layer linearly transforms an input hidden state , where with hidden dimension into the final output using the matrix product for each channel :
@@
In this paper, we implement a hypernetwork (3) as a MLP with one hidden layer that that takes a learnable embedding vector for each time series component (treating the number of channels as the batch size) and outputs the corresponding weights for the final prediction layer (2) of an arbitrary base forecasting model. In the simplest case of an MLP without hidden layers, the weights of the base model are obtained using a linear transform:
@@
The central contribution of our hypernetwork-based approach (HN-MVTS) lies in utilizing a hypernetwork to generate part of the parameters of a forecasting model in a computationally efficient yet scalable and data-adaptive manner. In traditional MVTS models, all output channels typically share the same output projection layer (CD), or each channel has its own separate model (CI). Our approach aims to combine both benefits by introducing a shared neural forecasting model, while tailoring its final layer to each channel via a learnable hypernetwork. The difference between CD, CI, and our HN-MVTS is shown in Fig. 1. The former is the most common model, with a single transform of input data into outputs. The CI approach learns separate models , and processes each channel independently. Our HN-MVTS saves the advantages of both types of models. If some components and of time series are similar, their embeddings will be close to each other , and, hence, the training data for the component will have more influence to learn the weights for the component, and vice versa. In the ideal case of clustered embeddings , a single (global) model (Montero-Manso and Hyndman 2021) is trained for the union of these components. In contrast, if the embeddings of and channels are significantly different, a CI mode is activated, in which the forecasting parameters are trained independently. This mechanism allows the model to adaptively interpolate between CD and CI behavior based on the learned embedding, without manual architecture changes or channel grouping.
@@
Some examples of adding our HN-MVTS to various architectures are shown in Fig. 2. The above-mentioned motivation of mixing CD and CI models is especially clear for a forecasting model with one linear layer (left part of this figure), where the columns of the weight matrix are learned based on the similarity between embeddings of time series components. However, our approach can be easily applied to MLP (Fig. 2b) or more complex architectures such as transformers (Fig. 2c), as only the weights of the last prediction layer are defined by the hypernetwork. It may be used for all layers, but the number of trained parameters will be significantly greater, making it more challenging to train the model. A practitioner can control the number of excess parameters by managing the complexity of the hypernetwork.
@@
In all experiments, we use the following straightforward implementation of HN-MVTS. First, we initialize embeddings from the Pearson product-moment correlation coefficients across channels (Nguyen-Thai et al. 2024), which are projected onto the principal components of dimension computed using the training-only splits. Our experiments indicate that alternative (random) initialization results in slightly higher MSE. Next, we attach a simple MLP (4) and learn its weights and embeddings from given data together with the base model (1) by minimizing MSE. Our simplest implementation (4) adds only parameters to learn (with addition of if embeddings are learnable). As a result, the number of parameters can be even less than that of the CI approach, which applies the base model separately with its parameters for each channel. As the number of learnable weights is reasonably small, the training time is only slightly higher than that of the base forecasting model.
Moreover, since the trained hypernetwork outputs fixed weights that do not depend on the specific input time series, we can compute them only once after the training procedure is complete. Hence, we remove a hypernetwork and copy weights (4) into the last linear layer of the base model (2). As a result, the inference time won’t be affected (though deployment overheads, e.g., I/O, memory layout, may still cause minor variation), making our HN-MVTS as fast as the base model.
@@
Thus, the proposed approach uses a learnable component embedding combined with a hypernetwork to produce channel-specific final-layer weights as a data-adaptive regularizer, integrates seamlessly with diverse backbone architectures, and preserves inference (Savchenko 2016). These points with empirical gains discussed in the next section are a substantive contribution beyond the straightforward reapplication of existing hypernetwork’s ideas.
@@
We evaluated the proposed HN-MVTS with contemporary multivariate time series forecasting models of different architectures: 1) Linear model: DLinear (Zeng et al. 2023), 2) MLP-based model: TSMixer (Chen et al. 2023), 3) Convolutional neural network-based model: ModernTCN (Luo and Wang 2024), and 4) Transformer-based models: PatchTST (Nie et al. 2023) and inverse Transformer (Liu et al. 2024b). The source code for all models was taken from the original repositories of their authors.
@@
We leverage eight publicly available open datasets that are conventionally used in the MVTS literature (Shao et al. 2025), covering energy, traffic, and weather forecasting applications: ECL (UCI Electricity Consuming Load) with electricity consumption of hundreds of points/clients; 15-minute-level datasets (ETTm1, ETTm2) from the Electricity Transformer Temperature database (Zhou et al. 2021) with one oil and six load features of electricity transformers from 2016 to 2018; Weather (Max-Planck-Institute Weather Dataset for Long-term Time Series Forecasting) with weather indicators recorded every 10 min for 2020, and Traffic Flow from Performance Measurement System (PEMS03, PEMS04, PEMS07, PEMS08) collected by California Transportation Agencies (CalTrans) every 30 seconds and aggregated to 5 minutes.
Each dataset exhibits diverse temporal and structural properties, allowing us to assess the robustness and generalizability of our method. Their statistics are summarized in Table 1. We split the time series into train, validation, and test sets using a 7:2:1 ratio. For the ETTm1 and ETTm2 datasets, following standard testing protocols, we use a split and take only the first 57600 timesteps. The validation set was used to obtain the best checkpoint of the model, and the results on the test set are reported.